# Separate the Wheat from the Chaff: A Post-Hoc Approach to Safety Re-Alignment for Fine-Tuned Language Models

### v1

@@ Abstract
Although large language models (LLMs) achieve effective safety alignment at the time of release, they still face various safety challenges. A key issue is that fine-tuning often compromises the safety alignment of LLMs. To address this issue, we propose a method named IRR (Identify, Remove, and Recalibrate for Safety Realignment) that performs safety realignment for LLMs. The core of IRR is to identify and remove unsafe delta parameters from the fine-tuned models, while recalibrating the retained ones. We evaluate the effectiveness of IRR across various datasets, including both full fine-tuning and LoRA methods. Our results demonstrate that IRR significantly enhances the safety performance of fine-tuned models on safety benchmarks, such as harmful queries and jailbreak attacks, while maintaining their performance on downstream tasks. The source code is available at: https://anonymous.4open.science/r/IRR-BD4F.
@@ 1 Introduction
In the "release and fine-tuning" paradigm, LLMs acquire delta parameters through fine-tuning, which enhance their performance on downstream tasks. However, this process often compromises the safety mechanisms established during safety alignment, reducing their value as reliable AI services. Specifically, training data that mixes harmful data with benign data, or consists entirely of benign data, can significantly compromise the safety alignment of LLMs (Bhardwaj and Poria, 2023; Wan et al., 2023; Zhan et al., 2024; Qi et al., 2023). Given the widespread adoption of the “release and fine-tuning” paradigm and its associated risks, a key objective is to ensure the safety realignment of fine-tuned models while maintaining their performance on downstream tasks.
@@
Figure 1: The illustration presents post-hoc approaches for safety realignment. Our method, IRR, first identifies and removes unsafe delta parameters, then recalibrates the remaining ones. The IRR approach enhances safety while maintaining performance.
@@
To achieve this objective, a straightforward method is to directly modify the parameters of fine-tuned models. A notable example is RESTA (Bhardwaj et al., 2024), which improves model safety by incorporating a safety vector into all parameters. However, current methods overlook the relationship between parameters associated with safety and those linked to downstream tasks in fine-tuned models (Frankle and Carbin, 2018; Panigrahi et al., 2023; Wei et al., 2024). As a result, applying the same safety modifications to all parameters without distinction may unintentionally harm those critical for downstream task performance. Inspired by recent work on removing redundant delta parameters from fine-tuned models (Panigrahi et al., 2023; Yu et al., 2024), we propose separating unsafe delta parameters from fine-tuned models. As shown in Figure 1, by removing unsafe delta parameters and recalibrating the retained parameters, we can achieve safety realignment while maintaining performance on downstream tasks.
@@
Based on these insights, we propose IRR (Identify, Remove, and Recalibrate for Safety Realignment), a simple yet effective post-hoc method for the safety realignment of fine-tuned models. Specifically, the IRR method involves three steps: (1) We use safety vectors to represent the parameter changes that move the model from an unsafe state to a safety-aligned state. Next, we identify delta parameters that interfere with parameters important for safety alignment and mark them as unsafe. These interferences arise due to sign disagreement with the safety vector. (2) We remove the unsafe delta parameters from the fine-tuned models. The remaining delta parameters are retained. (3) Unsafe delta parameters can compromise model safety but may be important for downstream task performance. Therefore, removing unsafe delta parameters can lead to performance degradation. To mitigate this, we recalibrate the retained delta parameters using precise weight compensation based on the inverse of the Hessian matrix, aiming to maintain performance.
@@
The main contributions of this work are summarized as follows:
  • We propose IRR, a novel safety realignment method that improve safety through three steps: identify, remove, and recalibrate.
@@ 2 Related Work
The safety of LLMs aims to mitigate potential safety risks arising from misuse or malicious use. Recent studies have identified vulnerabilities in the safety alignment of LLMs. Yang et al. (2023); Bhardwaj and Poria (2023); Zhan et al. (2024) demonstrated that even fine-tuning on small amounts of harmful data can significantly impact the safety alignment of LLMs. Qi et al. (2023) used more practical datasets, such as identity shift data and benign data like Alpaca, to undermine the safety alignment of LLMs.
To address the safety compromises introduced by fine-tuning, current methods focus on three main phases: (1) Pre-processing phase, where Zhao et al. (2023a) uses catastrophic forgetting to filter harmful data; (2) Fine-tuning phase, where Huang et al. (2024) limits parameter updates to reduce safety loss; (3) Post-hoc phase, where Bhardwaj et al. (2024) employs a merging safety vector approach to enhance model safety, and Zhao et al. (2024) introduces a patch to the safety vector to mitigate over-safety issues. Additionally, Hsu et al. (2024) proposes the Safe LoRA method, and Yi et al. (2024) utilizes sub-network search techniques to train a safety sub-network within the fine-tuned model. Inference-time safety has also gained attention, with Hazra et al. (2024) removing harmful vectors and adjusting the latent space, and Xu et al. (2024) improving content safety by increasing the probability of safe tokens during decoding.
@@
As neural network models grow in size, model pruning techniques have been widely adopted to reduce computational costs (Cheng et al., 2017; Liang et al., 2021). The goal of model pruning is to remove unnecessary parameters while maintaining model performance (Zhu and Gupta, 2017). The key to successful pruning lies in evaluating the importance of parameters and removing those with low importance scores to compress the model. For example, Liu et al. (2021) used the Fisher matrix to compute parameter importance and removed those with low scores. Frantar and Alistarh (2023) introduced SparseGPT, forming importance scores by solving a layer-wise reconstruction problem, while Sun et al. (2024) proposed the Wanda score, which assesses parameter importance using joint weight/activation metrics. Although our approach shares the concept of parameter removal with model pruning, our focus is on removing certain delta parameters rather than removing entire parameters.
@@ 3 Approach
We propose IRR, a novel method for safety realignment of fine-tuned models, which effectively restores safety while maintaining downstream task performance. The overall framework of IRR is illustrated in Figure 2, and consists of the following three steps: (1) Identify the Unsafe Delta Parameters. This step identifies delta parameters that interfere with important parameters for safety alignment and marks these interfering delta parameters as unsafe. (2) Remove the Unsafe Delta Parameters. These unsafe delta parameters are removed and reverted to their original safe pre-trained values, improving the safety of fine-tuned models. (3) Recalibrate the Retained Parameters. Since some unsafe delta parameters may significantly affect downstream task performance, removing them could degrade performance. To mitigate this, we compute compensatory values and add them to the retained parameters.
@@ 4 Experimental Setup
We conducted experiments using both full fine-tuning and LoRA (Hu et al., 2022). The results for full fine-tuning are presented in the main text, while LoRA results are in Appendix B.
@@
Our experiments are conducted on the widely used open-source model Llama-2-7b-chat (Touvron et al., 2023), which has been fine-tuned to follow instructions, align with human preferences, and ensure strong safety. Additionally, we perform LoRA fine-tuning experiments based on the Llama-3-8B-Instruct (Dubey et al., 2024). Supervised fine-tuning (SFT) is conducted using the LLaMA Factory  [^1] https://github.com/hiyouga/LLaMA-Factory, and the resulting models are referred to as domain-specific fine-tuned models.
@@
According to Bhardwaj et al. (2024), we define the safety vector $\delta_{safe}$ as the difference in parameters between the aligned and unaligned models. The unaligned model is fine-tuned using a harmful question-answer dataset. We extracted 1,000 labeled harmful question-answer pairs from the BeaverTails dataset (Ji et al., 2024) for training.
@@
Table 1: We evaluate safety on harmful benchmarks and jailbreak attacks. A higher safety score indicates better safety, and $\Delta$ represents the difference in safety compared to the original model. Higher performance in downstream tasks reflects better capability. The best and second-best results are highlighted in bold and underlined, respectively.
@@
For the harmful query benchmark, we utilized three datasets: 1) CATQA (Bhardwaj et al., 2024), a multilingual dataset encompassing English, Chinese, and Vietnamese; 2) HEx-PHI (Qi et al., 2023), which contains 330 harmful queries collected based on the usage policies of Meta and OpenAI; 3) Salad-Base (Li et al., 2024), covering 6 domains, 16 tasks, and 66 categories. We performed stratified sampling on 10% of the Salad-Base dataset and obtained a total of 2,100 harmful queries.
@@ 5.2 Ablation Study
We ablated the identification step (IRR w/o ID) and replaced it with a random selection of delta parameters to be removed. As shown in Figure 5, skipping the identification step typically results in a significant drop in downstream task performance as a trade-off for improved safety. This highlights the critical importance of identifying unsafe parameters.
@@
IRR uses safety interference together with the safety importance score to identify unsafe delta parameters. We ablated the safety interference strategy in the identification step (IRR w/o SI). As shown in Figure 5, relying only on the safety importance score to identify unsafe delta parameters also leads to significant degradation in downstream task performance.
@@ Appendix A Experimental Details
For the Math, Code, and Chinese in Table 1, we set varying values for $\rho$, specifically 60%, 10%, and 80%, respectively. However, for each task, we can optimize the hyperparameters to achieve the best trade-off between downstream task performance and safety, as illustrated in Figure 4. We employed greedy decoding as our generation strategy. All experiments were conducted on 4 $\times$ A100 80GB GPUs.
@@
Llama-2 training and inference are conducted with the default system prompt.
@@
Figure 8: We show the trend of “task performance vs. safety score” based on the Harmful Benchmark. Our method, IRR, outperforms baseline methods, maintaining strong downstream task performance as safety improves. Even when safety is close to that of the original model, downstream performance shows little degradation.
@@
Figure 9: We show the trend of “task performance vs. safety score” based on the Jailbreak Attack. Our method, IRR, outperforms baseline methods, maintaining strong downstream task performance as safety improves. Even when safety is close to that of the original model, downstream performance shows little degradation.
@@ Appendix D Ablation Experiments
Figure 11: We present the results of the IRR ablation study using “task performance vs. safety” curves. Our method effectively identifies unsafe delta parameters and, combined with the calibration step, successfully preserves downstream task performance.
@@ Appendix E Details about Llama-3-8B-Instruct experiment
In the Llama3 experiments, we fine-tuned Meta-Llama-3-8B-Instruct (Dubey et al., 2024) using LoRA. The query and value matrices in LoRA were tuned with a rank of $r=8$. The training batch size was set to 64, and the learning rate configured as $2.5\times 10^{-4}$. Fine-tuning was performed on the MathInstruct dataset (Yue et al., 2024).
@@
Meta-Llama-3-8B-Instruct does not include a default system prompt, so no system prompt is added during training or inference.
Appendix F Computational Complexity of IRR
@@
(1) Initial Hessian Calculation: The time complexity for calculating the initial Hessian matrix is $O(nd^{2}_{\text{col}})$, where $n$ represents the number of input samples and $d_{\text{col}}$ is the number of columns in the matrix.
(2) Hessian Inversion Iteration: The iterative inversion of the Hessian matrix has a time complexity of $O(d^{3}_{\text{col}})$, which remains manageable even for large models.
(3) Reconstruction Process: The pruning or reconstruction process based on the inverse Hessian involves a complexity of $O(d^{3}_{\text{col}}+d^{2}_{\text{row}}d_{\text{col}})$, where $d_{\text{row}}$ denotes the number of rows in the matrix. This ensures that the process is computationally feasible even for models with a large number of parameters.

### v2

@@ Abstract
Although large language models (LLMs) achieve effective safety alignment at the time of release, they still face various safety challenges. A key issue is that fine-tuning often compromises the safety alignment of LLMs. Previous post-hoc methods make safety modifications to the model parameters as a whole, which can result in performance degradation. We propose a method named IRR (Identify, Remove, and Recalibrate for Safety Realignment) that performs safety realignment for LLMs. The core of IRR is to identify and remove unsafe delta parameters from the fine-tuned models, while recalibrating the retained parameters. We evaluate the effectiveness of IRR across various datasets, including both full fine-tuning and LoRA methods. Our results demonstrate that IRR significantly enhances the safety performance of fine-tuned models, while maintaining their performance on downstream tasks. The source code is available at: https://github.com/pikepokenew/IRR.
@@ 1 Introduction
In the “release and fine-tuning” paradigm, the parameters of LLMs change, and these changes are known as delta parameters, which improve performance on downstream tasks. However, this process often compromises the safety mechanisms established during safety alignment, reducing their value as reliable AI services. Specifically, training data that mixes harmful data with benign data, or even consists entirely of benign data, can significantly compromise the safety alignment of LLMs (Bhardwaj and Poria, 2023; Qi et al., 2023; Wan et al., 2023; Zhan et al., 2024; Huang et al., 2024b). Given the widespread use of the “release and fine-tuning” paradigm and its associated risks, a key objective is to ensure the safety realignment of fine-tuned models while maintaining their performance on downstream tasks.
@@
Figure 1: The illustration presents post-hoc approaches for safety realignment. Our method, IRR, first identifies and removes unsafe delta parameters, then recalibrates the remaining ones.
@@
Recent works show directly modifying the parameters of the fine-tuned model can effectively and flexibly achieve this objective. A notable example is RESTA (Bhardwaj et al., 2024), which improves safety by merging a safety vector into all parameters (Figure 1). However, current methods overlook the safety relevance and task relevance of parameters (Frankle and Carbin, 2018; Panigrahi et al., 2023; Wei et al., 2024), as well as the trade-offs between them. As a result, applying the same safety modifications to all parameters without distinction may unintentionally harm those critical for downstream task performance. Additionally, some research shows that removing redundant delta parameters from fine-tuned models can restore the abilities that were forgotten due to fine-tuning (Panigrahi et al., 2023; He et al., 2024; Zhu et al., 2024). Inspired by this, we suggest separating unsafe delta parameters from the fine-tuned models to restore safety and recalibrating the retained parameters to minimize the performance loss, as shown in Figure 1. This method can achieve a better balance between safety and downstream task performance.
@@
Based on these insights, we propose IRR (Identify, Remove, and Recalibrate for Safety Realignment), a simple post-hoc method for the safety realignment of fine-tuned models. Specifically, the IRR method involves three steps: (1) To identify unsafe delta parameters in the fine-tuned model, we introduce a safety vector. This vector represents the changes that move the model from an unsafe state to a safe state. If the sign of a delta parameter disagrees with the safety vector, it can interfere with the model safety. We identify the delta parameters that cause safety interference and are important for safety as unsafe. (2) We remove the unsafe delta parameters from the fine-tuned models. (3) Unsafe delta parameters can compromise safety but may be important for downstream task performance. Removing them might reduce performance. To mitigate this, we recalibrate the retained parameters using precise weight compensation based on the inverse of the Hessian matrix.
@@
Our contributions can be summarized as follows:
  • We propose IRR, a novel safety realignment method that improves safety through three steps: identify, remove, and recalibrate.
@@ 2 Related Work
The safety of LLMs aims to mitigate potential safety risks arising from misuse or malicious use. Recent studies have identified vulnerabilities in the safety alignment of LLMs. Yang et al. (2023); Bhardwaj and Poria (2023); Zhan et al. (2024); Huang et al. (2024b) demonstrated that even fine-tuning on small amounts of harmful data can significantly impact the safety of LLMs. Qi et al. (2023) used more practical datasets, such as identity shift data and benign data like Alpaca, to undermine the safety alignment of LLMs.
To address the safety compromises introduced by fine-tuning, current methods focus on four main phases: (1) Pre-processing phase, where Zhao et al. (2023a) uses catastrophic forgetting to filter harmful data; (2) Fine-tuning phase, where Huang et al. (2024c) limits parameter updates to reduce safety loss; (3) Post-hoc phase after fine-tuning, where Bhardwaj et al. (2024) employs a merging safety vector approach to enhance safety, and Zhao et al. (2024) introduces a patch to the safety vector to mitigate over-safety issues. Additionally, Hsu et al. (2024) proposes the Safe LoRA, and Yi et al. (2024) trains a safety sub-network within the fine-tuned model. Huang et al. (2024a) employs model pruning to remove harmful parameters from the model. These methods do not require intervention during the inference time; (4) Inference time phase, where Hazra et al. (2024) removed harmful vectors and adjusted the latent space, while Xu et al. (2024) improved model safety by increasing the probability of safety tokens during the decoding.
@@
As neural network models grow in size, model pruning techniques have been widely adopted to reduce computational costs (Cheng et al., 2017; Liang et al., 2021). The goal of model pruning is to remove unnecessary parameters while maintaining model performance (Zhu and Gupta, 2017). The key to successful pruning is to assess the importance of parameters. For example, Liu et al. (2021) used the Fisher matrix to compute parameter importance and removed those. Frantar and Alistarh (2023) introduced SparseGPT, forming importance scores by solving a layer-wise reconstruction problem, while Sun et al. (2024) proposed the Wanda score, which assesses parameters using joint weight/activation metrics. Although our approach involves parameter removal similar to model pruning, we specifically focus on removing delta parameters rather than entire parameters.
@@ 3 Approach
We propose IRR, a novel method for safety realignment of fine-tuned models, which effectively restores safety while maintaining downstream task performance. The overall framework of IRR is illustrated in Figure 2, and consists of the following three steps: (1) Identify the Unsafe Delta Parameters. This step identifies delta parameters that interfere with important parameters for safety alignment and marks these interfering delta parameters as unsafe. (2) Remove the Unsafe Delta Parameters. These unsafe delta parameters are removed and reverted to their original safe pre-trained states, improving the safety of fine-tuned models. (3) Recalibrate the Retained Parameters. Since some unsafe delta parameters may significantly affect downstream task performance, removing them could degrade performance. To mitigate this, we compute compensatory values and add them to the retained parameters.
@@ 4 Experimental Setup
We conducted experiments using both full fine-tuning and LoRA (Hu et al., 2022). The results for full fine-tuning are presented in the main text, while LoRA results are in Appendix B. Experiments for more scenarios can be found in Appendix F and G.
@@
Our experiments are conducted on the widely used open-source model Llama-2-7b-chat (Touvron et al., 2023), which has been fine-tuned to follow instructions, align with human preferences, and ensure strong safety. Additionally, we perform LoRA fine-tuning experiments based on the Llama-3-8B-Instruct (Dubey et al., 2024). Supervised fine-tuning (SFT) is conducted using the LLaMA Factory  [^1] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024), and the resulting models are referred to as domain-specific fine-tuned models.
@@
According to Bhardwaj et al. (2024), we define the safety vector $\delta_{safe}$ as the difference in parameters between the aligned and unaligned models. The unaligned model is fine-tuned using a harmful question-answer dataset. We extracted 1,000 labeled harmful question-answer pairs from the BeaverTails dataset (Ji et al., 2024) for training. Safe LoRA and RESTA use the same safety vector.
@@
Table 1: We evaluate safety using harmful benchmarks and jailbreak attacks. A higher safety score indicates better safety, while $\Delta$ represents the difference in safety score compared to the original model. Higher performance in downstream tasks reflects better capability. The best and second-best results are highlighted in bold and underlined.
@@
For the harmful query benchmark, we used three datasets: 1) CATQA (Bhardwaj et al., 2024), a multilingual dataset with English, Chinese, and Vietnamese; 2) HEx-PHI (Qi et al., 2023), which includes 330 harmful queries based on usage policies from Meta and OpenAI; 3) Salad-Base (Li et al., 2024), covering 6 domains, 16 tasks, and 66 categories. We did stratified sampling on 10% of the Salad-Base, resulting in 2,132 harmful queries.
@@ 5.2 Ablation Study
We ablated the identification step (IRR w/o ID) and replaced it with a random selection of delta parameters to be removed. As shown in Figure 5, skipping the identification step typically results in a significant drop in downstream task performance as a trade-off for improved safety. This highlights the critical importance of identifying unsafe delta parameters.
@@
IRR uses safety interference together with the safety importance score to identify unsafe delta parameters. We ablated the safety interference strategy in the identification step (IRR w/o SI). As shown in Figure 5, relying only on the safety importance score to identify unsafe delta parameters also leads to significant degradation in downstream task performance. This indicates that safety interference plays an important role.
@@ 7 Limitations
Our IRR.
Given the complex architecture of LLMs, our approach for obtaining safety vectors and evaluating the safety importance scores of parameters in the IRR method is relatively simple. Developing more robust and precise methods for these steps is necessary and should be a focus of future investigations.
@@
8 Ethical consideration
Ensuring ethical applications of artificial intelligence is critical. Our safety realignment method IRR enhances the safety of fine-tuned language models by reducing harmful content. The identification and removal operations effectively reduce harmful responses in fine-tuned models, while calibration ensures strong downstream task performance. Our framework demonstrates its effectiveness in improving the safety of fine-tuned models across different datasets. We advocate for ongoing collaboration among researchers, policymakers, and industry stakeholders to ensure that artificial intelligence development prioritizes human values, fairness, and safety. We remain committed to continuously evaluating and improving our approach to address ethical challenges.
@@ Appendix A Experimental Details
For the Math, Code, and Chinese in Table 1, we set varying values for $\rho$, specifically 60%, 10%, and 80%, respectively. However, for each task, we can optimize the hyperparameters to achieve the best trade-off between downstream task performance and safety, as illustrated in Figure 3 and Figure 4. To plot the trade-off between safety and downstream task performance, the mask ratio starts at 10% and increases in increments of 10% until it reaches 100%. The calculation of compensatory values in the recalibration step requires downstream task data. For this purpose, we extracted 1,000 samples from the corresponding downstream task dataset. We employed greedy decoding as our generation strategy. All experiments were conducted on 4 $\times$ A100 80GB GPUs.
@@
Both the training and inference of Llama-2 use the default system prompt and the chat template.
@@
Figure 8: LoRA experimental results. We present the trends of "task performance versus safety score" on the harmful benchmark. Our method, IRR, outperforms baseline methods by improving safety while maintaining strong downstream task performance.
@@
Figure 9: LoRA experimental results. We present the trends of "task performance versus safety score" on the Jailbreak attacks. Our method, IRR, outperforms baseline methods by improving safety while maintaining strong downstream task performance.
@@ Appendix D Ablation Experiments
Figure 11: We present the results of the IRR ablation study using “task performance vs. safety” curves on harmful benchmarks. Our method effectively identifies unsafe delta parameters and, combined with the calibration step, successfully preserves downstream task performance.
@@ Appendix E Details about Llama-3-8B-Instruct experiment
In the Llama3 experiments, we fine-tuned Meta-Llama-3-8B-Instruct (Dubey et al., 2024) using LoRA. The query and value matrices in LoRA were tuned with a rank of $r=8$. The training batch size was set to 64, and the learning rate configured as $2.5\times 10^{-4}$. Fine-tuning was performed on the MathInstruct dataset (Yue et al., 2024). We set the mask ratio in the IRR method to 1%, 2%, 3%, 4%, and 5%.
@@
Meta-Llama-3-8B-Instruct does not include a default system prompt, so no system prompt is added during training or inference. Both the training and inference of Llama-3 use the chat template.
Figure 13: Results of IRR and RESTA using the new safety vector.
Appendix F IRR with Different Safety Vectors
In the main text, the IRR method uses the same safety vector as Safe LoRA and RESTA. To explore the effectiveness of the IRR method with different safety vectors, we sampled an additional 1,000 harmful question-answer pairs from Beavertails (Ji et al., 2024) to train a new safety vector using LoRA. We present the results of the IRR and RESTA methods using a new safety vector on the model fine-tuned on GSM8K (Cobbe et al., 2021) with LoRA.
As shown in Figure 13, the IRR maintains better safety and downstream task performance compared to the RESTA when using the new safety vector. This indicates that the choice of safety vector does not affect the effectiveness of the IRR.
Appendix G IRR against Malicious Fine-tuning
To further investigate the effectiveness of IRR, we examined the scenario of mixing malicious instruction data into the fine-tuning dataset. We sampled 1,000 harmful question-answer pairs from Beavertails (Ji et al., 2024) and mixed them with the GSM8K (Cobbe et al., 2021) dataset for fine-tuning, keeping other settings the same.
We have added a special configuration to the IRR method called IRRmore. This configuration allows for the removal of other delta parameters in ascending order of their safety importance scores, after all delta parameters that cause safety interference have been removed.
Figure 14: Results of IRR against malicious fine-tuning.
As shown in Figure 14, in the case of malicious fine-tuning, the IRR method still effectively enhances safety while maintaining performance on downstream tasks compared to the RESTA method. Additionally, the IRR method can be combined with data processing methods (Zhao et al., 2023a; Yang et al., 2024) before fine-tuning and parameter update constraints (Huang et al., 2024c; Qi et al., 2024) applied during fine-tuning to achieve better safety improvements.
Appendix H Computational Complexity of IRR
@@
Initial Hessian Calculation. The time complexity for calculating the initial Hessian matrix is $O(nd^{2}_{\text{col}})$, where $n$ represents the number of input samples and $d_{\text{col}}$ is the number of columns in the matrix.
Hessian Inversion Iteration. The iterative inversion of the Hessian matrix has a time complexity of $O(d^{3}_{\text{col}})$, which remains manageable even for large models.
Reconstruction Process. The pruning or reconstruction process based on the inverse Hessian involves a complexity of $O(d^{3}_{\text{col}}+d^{2}_{\text{row}}d_{\text{col}})$, where $d_{\text{row}}$ denotes the number of rows in the matrix. This ensures that the process is computationally feasible even for models with a large number of parameters.
@@
Data Size Time (s) Acc (%) Safety Score (%) Jailbreak Safety Score (%)
8 227.03 42.43 99.26 85.52
64 260.29 42.02 99.12 83.74
128 298.23 42.14 99.20 83.41
1,000 839.22 42.06 99.16 82.45
Table 2: Results of running IRR on the GSM8K full fine-tuned models
Data Size Time (s) Acc (%) Safety Score (%) Jailbreak Safety Score (%)
8 57.10 35.81 99.51 81.60
64 67.57 35.66 99.53 81.13
128 79.34 35.79 99.52 81.01
1,000 242.97 35.53 99.52 80.29
Table 3: Results of running IRR on the GSM8K LoRA fine-tuned models
Model Size Time (s) / Layer
7B 8.13
13B 10.12
70B 31.08
Table 4: Results of running IRR on the GSM8K LoRA fine-tuned models
Appendix I Time Consumption of IRR
To further investigate the time efficiency of the IRR method, we analyzed the time required to run IRR. Specifically, we measured the time needed to execute IRR on both full and LoRA fine-tuned models and estimated the time required when scaling to larger models. The testing platform used was an A100 80GB GPU.
I.1 Time for the IRR Method
We used the GSM8K (Cobbe et al., 2021) dataset to measure the time required for the IRR method on full and LoRA fine-tuned Llama-2-7b-chat models (Touvron et al., 2023). The IRR method requires a specific amount of downstream task data to compute the Hessian matrix. We used 8, 64, 128, and 1000 samples from downstream tasks. The mask ratio started at 10% and increased incrementally by 10% up to 100%. We reported the average results for time consumption, downstream task accuracy, safety scores, and jailbreak safety scores.
As shown in Table 2 and Table 3, using a smaller amount of calibration data can reduce the time required for the IRR method without significantly affecting downstream task performance or safety scores. This demonstrates that the IRR method can achieve faster safety realignment with less calibration data. For example, when utilizing the IRR method on a LoRA fine-tuned 7B model with 64 calibration samples, only 67.57 seconds are needed. In comparison, fine-tuning the LoRA model on the GSM8K dataset for 5 epochs takes at least 1.55 hours. Furthermore, the IRR method allows for safety realignment without increasing inference time, highlighting its advantages.
I.2 Estimating Time Efficiency for Scaling to Larger Models
Next, we will estimate the running time of the IRR method when it is applied to larger models. We used 64 samples from downstream tasks to calculate the Hessian matrix and measured the time needed for the IRR method to run on a single layer of the model.
According to the Table 4, for larger models, such as the full fine-tuned 40-layer Llama2-13B, the IRR method takes about 404.8 seconds to complete. For the 80-layer 70B model, it takes approximately 2,486.4 seconds.
Appendix J Limited GPU Memory Resources
In scenarios with limited GPU memory resources, the IRR method, based on SparseGPT, computes the inverse Hessian matrix by traversing each layer of the model. After calculating the Hessian matrix for the current layer, it proceeds to the next layer’s computation. This approach theoretically allows for only the current layer to be loaded into GPU memory during the calculations, thereby enabling safe realignment within the constraints of limited GPU memory resources.